OpenStudy (anonymous):

sin(a)-icos(a)?=cos(a-pi/2)+isin(a-pi/2)

OpenStudy (anonymous):

\[\cos\left(a-\frac{\pi}{2}\right)=\cos a\cos\frac{\pi}{2}+\sin a\sin\frac{\pi}{2}=\sin a\] Likewise, \[\sin\left(a-\frac{\pi}{2}\right)=\sin a\cos\frac{\pi}{2}-\sin\frac{\pi}{2}\cos a=-\cos a\] which establishes the identity.